OpenStudy (anonymous):

Find the angle,theta between vectors; round to two decimal places: A) u=<3,2>, v=<4,0> B) u= 2i-3j, v=i-2j

ganeshie8 (ganeshie8):

familiar wid "dot product" ?

OpenStudy (anonymous):

kind of

ganeshie8 (ganeshie8):

\(\large \vec{a} \bullet \vec{b} = |\vec{a} | ~|\vec{b}|~ \cos(\color{red}{\theta})\)

ganeshie8 (ganeshie8):

there is the angle.. take dot product and solve it

ganeshie8 (ganeshie8):

A) u=<3,2>, v=<4,0> first find below : \(u.v = ?\) \(|u| = ?\) \(|v| = ?\)

OpenStudy (anonymous):

3,2? 4,0?

ganeshie8 (ganeshie8):

yes, \(u . v = <3, 2> . <4, 0> = 3*4 + 2*0 = 12\)

ganeshie8 (ganeshie8):

\(|u| = \sqrt{3^2 + 2^2} = \sqrt{13}\) \(|v| = \sqrt{4^2 + 0^2} = 4\)

ganeshie8 (ganeshie8):

plug these 3 values in the dot product formula

ganeshie8 (ganeshie8):

\(\large \vec{a} \bullet \vec{b} = |\vec{a} | ~|\vec{b}|~ \cos(\color{red}{\theta})\) \(\large 12 = \sqrt{13} \times 4~ \cos(\color{red}{\theta})\)

ganeshie8 (ganeshie8):

see if u can solve \(\color{red}{\theta}\)

OpenStudy (anonymous):

I cant

ganeshie8 (ganeshie8):

okay, its easy.. u need to use calucaltor : \(\large 12 = \sqrt{13} \times 4~ \cos(\color{red}{\theta})\) \(\large \frac{3}{\sqrt{13}} = \cos (\color{red}{\theta})\)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yea, im following you

ganeshie8 (ganeshie8):

\(\large \frac{3}{\sqrt{13}} = \cos (\color{red}{\theta})\) take \(\cos^{-1}\) both sides \(\large \cos^{-1}( \frac{3}{\sqrt{13}}) = \color{red}{\theta} \)

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

rounding to two decimal places 0.59?

ganeshie8 (ganeshie8):

give them angle in DEGREES..

ganeshie8 (ganeshie8):

0.59 is the angle in radians... but usually ur teachers expect u to give angle in DEGREEs..

OpenStudy (anonymous):

okay, so 33.69 degrees

ganeshie8 (ganeshie8):

Correct !!

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

and I do b the same way

ganeshie8 (ganeshie8):

yes, same way : find below 3 things : 1) u.v 2) |u| 3) |v|

ganeshie8 (ganeshie8):

and plug them in "dot product" formula

ganeshie8 (ganeshie8):

u wana give it a try ? :)

ganeshie8 (ganeshie8):

B) u= 2i-3j, v=i-2j

OpenStudy (anonymous):

yea, give me a few minutes

ganeshie8 (ganeshie8):

take ur time

OpenStudy (anonymous):

okay so I have 8=√13*√5 cos Θ

OpenStudy (anonymous):

im a little lost on what's next

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

8=√13*√5 cos Θ

ganeshie8 (ganeshie8):

isolate cos Θ first

OpenStudy (anonymous):

so I would divide by √5?

ganeshie8 (ganeshie8):

divide by √13*√5

OpenStudy (anonymous):

so i would put in my calculator √13/√5?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

8=√13*√5 cos Θ divide √13*√5 both sides 8/(√13*√5 ) = cos Θ

ganeshie8 (ganeshie8):

take \(\cos^{−1}\) both sides to get Θ

OpenStudy (anonymous):

oh okay. I think im getting the hang of it. You are very helpful. Thank you again.

ganeshie8 (ganeshie8):

np :) rounded to two decimal places teh angle wud be \(7.13\) degrees